OpenStudy (chupacabraj):

Finding local max/min of a multivariable function. I get stuck in the algebra part.

OpenStudy (chupacabraj):

\[F(x,y)= (x-y)(1-xy)\] FOIL \[f(x,y) = x-x^2y-y+xy^2\]

OpenStudy (chupacabraj):

\[f_{xx}= 1-2xy+y^2\] \[f_{yy}= -x^2-1+x\]

OpenStudy (chupacabraj):

So I get stuck when trying to find the critical points. I attempted by solving by elimination but it turned out ugly.

zepdrix (zepdrix):

If this is f,\[\rm f(x,y) = x-x^2y-y+xy^2\]then don't you mean to call this f_x? The first derivative?\[\rm f_{x}= 1-2xy+y^2\]

zepdrix (zepdrix):

\[\rm f_{xx}=-2y\]

OpenStudy (owen3):

Did you solve \( \nabla f (x,y) = \vec 0 \)

OpenStudy (owen3):

$$ -2xy+y^2+1 = 0 \\ -x^2+2xy-1 = 0 \\ \, \\ \\ x=1, ~ y = 1 \\ x = -1 , ~ y = -1 $$

OpenStudy (chupacabraj):

Oh that's right, I meant first derivatives.

OpenStudy (chupacabraj):

How did you solve for x and y?

OpenStudy (irishboy123):